# Decentralized Federated Anomaly Detection in Smart Grids: A P2P Gossip Approach

### v1

@@ III-D Gossip-based DFL Framework
            Calculate local gradient $g^{t}_{k}=\nabla f_{k}(W^{t})$ using $D_{k}$. Aggregate local gradient $g^{t}_{k}$ with previous peer, $g^{t}_{k-1}$ such that $\sum_{i=k-1}^{k}g^{t}_{agg}$.. Calculate normalized gradient $\bar{g}^{t}_{norm}=g^{t}_{agg}-\mu^{t}_{k}$. Calculate quantized gradient $\hat{g}^{t}_{k}=dpsign$ $(\bar{g}^{t}_{norm})$. Communicate $\hat{g}^{t}_{k}$ to random subset of nodes. end
@@
In order to enable the federated discrimination of cyber attack events from non-anomalous natural power system disturbances, we propose to implement two types of gossip protocols:
  1. 1.
    Random Gossip: This proposed gossip approach randomly selects client nodes within a network for communication. During each communication round, each $k$ computes its local gradient $\nabla f_{k}(Z^{t})$ using its training dataset $D_{k}$ and aggregates this gradient with that received from a randomly chosen peer node. This aggregated gradient is then normalized and quantized before being transmitted to another randomly selected peer node. This process repeats iteratively, facilitating decentralized communication and collaboration among nodes.
  2. 2.
    Epidemic Gossip: Epidemic approach, on the other hand, operates on the principle of spreading information to all client nodes within a network. Each $k$ communicates with a fixed subset of its neighboring $k$, known as its Epidemic Radius. Similar to random gossip, nodes compute, aggregate, normalize, and quantize gradients before transmitting them to neighboring nodes. However, unlike random gossip, Epidemic gossip ensures that each node eventually receives the information from all other nodes in the network and is implemented for contrastive empirical purposes.
By leveraging random gossip, nodes can efficiently exchange information without requiring explicit knowledge of the network topology. This enhances scalability and adaptability to changes in network conditions. On the other hand, Epidemic gossip ensures that all nodes eventually receive the information, thereby promoting information dissemination and consensus among nodes. As discussed in Section IV, we will compare and contrast the two implemented protocols to identify the superior one.
IV Empirical Evaluation & Discussions
In this section, we initially discuss the dataset used for the experimental set-up followed by comprehensive discussion of the empirical results, which validate the anomaly detection performance and robustness of our proposed solution as presented in Section III.
IV-A Dataset Description & Preparation
We conduct an empirical evaluation of our proposed federated decentralized anomaly detection framework for power systems using the openly available Mississippi State University and Oak Ridge National Laboratory Power System Attack (MSU-ORNL PSA) datasets, which are derived from industrial control systems. These datasets consist of 15 distinct data files, each representing various power system event scenarios. The binary dataset scenarios encompass both natural events, such as power system disturbances, and deliberate attack events, including cyberattacks. Natural events are categorized based on fault severity across distributed grids, ranging from 10-19%, 20-79%, to 80-90%. Conversely, attack events include: 1) remote tripping command injection, involving sending commands to IEDs to manipulate circuit breakers operations, 2) changes in IED settings to disrupt normal operations, and 3) false data injection attacks, intelligently altering power system values across parameters like reference power, voltage and current to mislead system operators. In the case of the three-class dataset, no events indicates normal operations load changes. The datasets consist of 128 features, encompassing measurements e.g. frequency, voltage phase angle and current magnitude, amongst others.
@@ IV-A Dataset Description & Preparation
IV-B Anomaly Detection Performance
@@
TABLE I: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in a Random Walk Gossip scenario using binary class dataset.
Performance Metric Learning Model Proposed Approach CNN LSTM RNN RBM
Anomaly Detection Accuracy 0.933 0.924 0.828 0.768 0.787
Anomaly Detection Precision 0.915 0.919 0.765 0.740 0.803
Anomaly Detection Recall 0.945 0.916 0.759 0.710 0.760
Anomaly Detection F-Score 0.939 0.921 0.773 0.742 0.764
TABLE II: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in a Random Walk Gossip scenario using the three-class dataset.
@@
TABLE III: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in an Epidemic scenario using the binary-class dataset.
Performance Metric Learning Model Proposed Approach CNN LSTM RNN RBM
Anomaly Detection Accuracy 0.922 0.904 0.814 0.752 0.792
Anomaly Detection Precision 0.908 0.899 0.788 0.710 0.774
Anomaly Detection Recall 0.917 0.903 0.754 0.715 0.761
Anomaly Detection F-Score 0.923 0.905 0.767 0.731 0.781
TABLE IV: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in an Epidemic Gossip scenario using the three-class dataset. Our proposed TAE model consistently outperforms all other models in terms of anomaly detection performance.
Figure 4: Comparison of confusion matrices for Epidemic and Random Walk (proposed) pertaining to a single binary-class dataset file evaluated using our proposed DFL approach. Upon analyzing the occurrences of false positives and false negatives, it becomes evident that Random Walk gossip DFL outperforms Epidemic gossip DFL approach.
Figure 5: Comparison of confusion matrices for Epidemic and Random Walk (proposed) pertaining to a single three-class dataset file evaluated using our proposed DFL approach. Upon analyzing the occurrences of false positives and false negatives, it becomes evident that Random Walk gossip DFL outperforms Epidemic gossip DFL approach.
@@
Table I vs. Table III & Table II vs. Table IV demonstrate the performance metrics of the two gossip-based decentralized federated learning approaches applied to anomaly detection in power systems, with Tables I & II representing the results obtained using the random walk gossip-based method Tables III & IV corresponding to the epidemic gossip-based approach. It is clear from comparing results of the tables that the our proposed random walk gossip-based approach outperforms the epidemic gossip-based method consistently in terms of all performance criteria. The random walk gossip-based approach in DFL encourages efficient communication and information exchange among grid regions by allowing each client node to randomly select its peer. This randomness encourages a diverse exchange of model updates,thus enhancing convergence and accuracy of model updates. Furthermore, a balanced distribution of updates is achieved, thus mitigating information bottlenecks and ensuring all grid regions contribute effectively to the learning process. Similarly, based on Fig. 4 & Fig. 5 whereby the confusion matrices depict the performance of the Random Walk gossip-based (above) and Epidemic gossip-based (below) approaches, it is evident based on analysis of the true positive and true negative values that the proposed Random Walk gossip-based DFL approach exhibits superior performance. The higher true positive and true negative rates in the Random Walk gossip-based DFL matrix suggest that it more accurately identifies and correctly classifies both positive and negative instances, making it a preferable choice for our particular anomaly detection scheme. Furthermore, in Fig. 6, we contrast the anomaly detection performance of our proposed Random Walk gossip based protocol against that of epidemic and conventional FL. Similarly, Random Walk outperforms Epidemic protocols in terms of misbehaviour detection. The conventional FL approach demonstrates the lowest performance due to its inherent limitations in facilitating efficient communication and information exchange among the distributed nodes.
IV-C Computational Overhead Analysis
@@
Following the successful validations of the anomaly detection performance of our proposed approach, we further assess the the computational overhead achieved through elimination of the centralized aggregator and gradient quantization as presented in Section III. Specifically, we contrast the time taken to train the neural networks in three distinct scenarios namely Random Walk gossip-based DFL, Epidemic gossip-based DFL and conventional FL. To ensure fairness during comparison, we utilize similar models with similar hyperparameters in all three setups. Fig. 7 illustrated a comparison of the computational efficiency among the two aforementioned types of gossip DFL protocols and a conventional FL scenario. The results reveal a consistent decrease in training time across all models when our proposed DFL-gradient quantization approach is employed against a conventional FL scheme. Notably, the Random Walk and Epidemic Gossip based FL approach demonstrate higher training time compared to the traditional FL framework which is mainly due to increased complexity and distributed aggregation. Furthermore, it is evident that Random Walk-based DFL demonstrates around 30% reduction in training time as compared to Epidemic-based DFL, indicating that Random Walk is less computationally demanding and much better suited for resource-constrained environments such as in the case of power systems.
IV-D Summary of findings
@@
In this section, we investigated the detection performance of our novel decentralized anomaly detection solution within power systems, focusing on the comparison between decentralized federated learning (DFL) and conventional federated learning (FL) methodologies. Through extensive experimentation and analysis, we found that our proposed Random Walk gossip algorithm consistently outperformed its epidemic gossip-based DFL and conventional FL counterparts. Specifically, the novel solution exhibited superior detection accuracy and more efficient communication and information exchange among participating grid regions. The decentralized nature of DFL enables more dynamic and adaptive learning processes, leveraging the collective intelligence of distributed grid regions while mitigating the limitations associated with centralized coordination mechanisms in conventional FL setups. Additionally, we investigated the computational overhead and found that the our proposed DFL method incurred negligibly higher time costs compared to conventional FL approaches which is due to the complexity of the aggregation process and could be further improved through efficient gradient quantization. However, random walk protocol requires 35% less training time further emphasizing its computational efficiency and scalability for anomaly detection for Smart Grids.
V Conclusion
@@
In this paper, we investigated the effectiveness of a Decentralized Federated Learning scheme for anomaly detection in the domain of power systems. To the best of our knowledge, we are the first to introduce a gossip-based solution to decentralized federated anomaly detection in the SG ecosystem. First, we proposed a Random Walk gossip-based anomaly detection that enables the transition from a conventional Federated Learning paradigm to a decentralized one while eliminating the need for a centralized aggregator. Secondly, we proposed a hybrid anomaly detection neural network namely a Transformer Autoencoder (TAE) model. Extensive experimental validations using a publicly available datasets show that the TAE achieves superior accuracy as opposed to other state-of-the-art benchmark neural networks. Lastly, we introduce a sign-based gradient quantization scheme to optimize communication efficiency within the DFL framework, resulting in a significant reduction in training time. Overall, the findings concluded that there is a compelling potential for developing anomaly detection solutions using DFL and that Random Walk has shone in terms of both detection performance and computational overhead.

### v2

@@ 3.4 Gossip-based DFL Framework
            Calculate local gradient $g^{t}_{k}=\nabla f_{k}(W^{t})$ using $D_{k}$. Aggregate local gradient $g^{t}_{k}$ with previous peer, $g^{t}_{k-1}$ such that $\sum_{i=k-1}^{k}g^{t}_{agg}$.. Calculate normalized gradient $\bar{g}^{t}_{norm}=g^{t}_{agg}-\mu^{t}_{k}$. Calculate quantized gradient $\hat{g}^{t}_{k}=dpsign$ $(\bar{g}^{t}_{norm})$. Communicate $\hat{g}^{t}_{k}$ to peer node. end
@@
4 Empirical Evaluation & Discussions
In this section, we initially discuss the dataset used for the experimental set-up followed by comprehensive discussion of the empirical results which validate the anomaly detection performance and robustness of our proposed solution as presented in Section 3.
4.1 Dataset Description & Preparation
We conduct an empirical evaluation of our proposed federated decentralized anomaly detection framework for power systems using the openly available Mississippi State University and Oak Ridge National Laboratory Power System Attack (MSU-ORNL PSA) dataset, which is derived from industrial control systems [16]. This dataset consists of 15 distinct data files, each representing various power system event scenarios. These scenarios encompass both natural events, such as power system disturbances, and deliberate attack events, including cyberattacks. Natural events are categorized based on fault severity across distributed grids, ranging from 10-19%, 20-79%, to 80-90%. Conversely, attack events include: 1) remote tripping command injection, involving sending commands to IEDs to manipulate circuit breakers; 2) changes in IED settings to disrupt normal operations; and 3) false data injection attacks, intelligently altering power system values across parameters like voltage and current to mislead system operators. The datasets consist of 128 features, encompassing measurements e.g. frequency, voltage phase angle and current magnitude, amongst others.
@@ 4.1 Dataset Description & Preparation
4.2 Anomaly Detection Performance
@@
Table 1: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in a Random Walk Gossip scenario. Our proposed TAE model consistently outperforms all other models in terms of anomaly detection performance.
@@
Table 2: Comparison of average anomaly detection performance of our proposed TAE model against other neural networks in an Epidemic scenario. Our proposed TAE model consistently outperforms all other models in terms of anomaly detection performance.
Figure 4: Comparison of confusion matrices for Epidemic (above) and RandomWalk (below) pertaining to a single data file evaluated using our proposed DFL approach. Upon analyzing the occurrences of false positives and false negatives, it becomes evident that Random Walk gossip DFL outperforms Epidemic gossip DFL approach.
@@
Table 1 and Table 2 demonstrate the performance metrics of the two gossip-based decentralized federated learning approaches applied to anomaly detection in power systems, with Table 1 representing the results obtained using the random walk gossip-based method Table 2 corresponding to the epidemic gossip-based approach. It is clear from comparing the two tables’ results that the random walk gossip-based approach outperforms the epidemic gossip-based method consistently in terms of all performance criteria. The random walk gossip-based approach in DFL encourages efficient communication and information exchange among grid regions by allowing each client node to randomly select its peer. This randomness encourages a diverse exchange of model updates,thus enhancing convergence and accuracy of model updates. Furthermore, a balanced distribution of updates is achieved, thus mitigating information bottlenecks and ensuring all grid regions contribute effectively to the learning process. Similarly, based on Fig. 4 whereby the confusion matrices depict the performance of the Random Walk gossip-based (above) and Epidemic gossip-based (below) approaches, it is evident based on analysis of the true positive and true negative values that Random Walk gossip-based DFL approach exhibits superior performance. The higher true positive and true negative rates in the Random Walk gossip-based DFL matrix suggest that it more accurately identifies and correctly classifies both positive and negative instances, making it a preferable choice for our particular scenario. Furthermore, in Fig. 5, we contrast the anomaly detection performance of the two gossip based protocols against that of conventional FL. The conventional FL approach demonstrates the lowest performance due to its inherent limitations in facilitating efficient communication and information exchange among the distributed nodes.
4.3 Computational Overhead Analysis
@@
Following the successful validations of the anomaly detection performance of our proposed approach, we further assess the the computational overhead achieved through elimination of the centralized aggregator and gradient quantization as presented in Section 3. Specifically, we contrast the time taken to train the neural networks in three distinct scenarios namely Random Walk gossip-based DFL, Epidemic gossip-based DFL and conventional FL. To ensure fairness during comparison, we utilize similar models with similar hyperparameters in all three setups. Fig. 6 illustrated a comparison of the computational efficiency among the two aforementioned types of gossip DFL protocols and a conventional FL scenario. The results reveal a consistent decrease in training time across all models when our proposed DFL-gradient quantization approach is employed against a conventional FL scheme. Notably, the Random Walk and Epidemic Gossip based FL approach demonstrate approximately 35% less training time compared to the traditional FL framework. Furthermore, it is evident that Random Walk-based DFL demonstrates around 3% to 5% reduction in training time as compared to Epidemic-based DFL, indicating that Random Walk is less computationally demanding and much better suited for resource-constrained environments such as in the case of power systems.
4.4 Summary of findings
@@
In this section, we investigated the detection performance of our novel decentralized anomaly detection solution within power systems, focusing on the comparison between decentralized federated learning (DFL) and conventional federated learning (FL) methodologies. Through extensive experimentation and analysis, we found that DFL methods by employing gossip-based communication protocols such as random walk and epidemic gossip, consistently outperformed the conventional FL scheme. Specifically, DFL methods exhibited superior detection accuracy and more efficient communication and information exchange among participating grid regions. The decentralized nature of DFL enables more dynamic and adaptive learning processes, leveraging the collective intelligence of distributed grid regions while mitigating the limitations associated with centralized coordination mechanisms in conventional FL setups. Additionally, we investigated the computational overhead and found that the conventional FL method incurred higher time costs compared to DFL approaches, further emphasizing the computational efficiency and scalability of decentralized federated anomaly detection for SGs.
5 Conclusion
@@
Throughout this manuscript, we investigated the effectiveness of a Decentralized Federated Learning (DFL) scheme for anomaly detection in the domain of power systems. To the best of our knowledge, we are the first to introduce a gossip-based solution to decentralized federated anomaly detection in the SG ecosystem. First, we proposed two gossip protocols namely Random Walk and Epidemic that enables the transition from a conventional Federated Learning paradigm to a decentralized one while eliminating the need for a centralized aggregator. Secondly, we proposed a hybrid anomaly detection neural network namely a Transformer Autoencoder (TAE) model. Extensive experimental validations using a publicly available dataset show that the TAE achieves superior accuracy as opposed to other state-of-the-art benchmark neural networks. Lastly, we introduce a sign-based gradient quantization scheme to optimize communication efficiency within the DFL framework, resulting in a significant reduction in training time. Overall, the findings concluded that there is a compelling potential for developing anomaly detection solutions using DFl and that Random Walk has shone in terms of both detection performance and computational overhead.